Please answer correctly or don’t answer at all thank you
have signhicant gold holdings. But many of these countries also have massive debes The following dats show the total value of product for nine countries (WondPress and Trading Economics websites, February especialily those in Europe, Many countrie 24, 2012) Cick on the datafile loge to reference the data DATA Gold Country 17.7 Prance Germany Indonesia 132 692 62 .Select for the total value of a countrys gold holdings( bitions) as the independent variable B. at
media%2Fe7d%2Fe7dfd259-221b-476a-b66c-1c
media%2Fe08%2Fe08f7f16-3e4e-4eb9-ad48-a1
media%2F8d3%2F8d3facd1-06c4-4230-a00e-fa
media%2F3f7%2F3f7594f9-0832-4ce2-9c90-98
Answer #1

XY 63 17.7 1115.13969 313.29 146 81.7 11928.2 21316 6674.89 203 83.2 16889.6 41209 6922.24 33 69.2 2283.6 1089 4788.64 147 119 17493 2160914161 36 63.7 2293.21296 4057.69 495 2500 98.01 3025 487 93.2 45388.4 237169 8686.24 50 9.9 62 553410 3844 ? ? = 1227.0. ? ? = 592.6 ? ?? = 101296.1. ? xx = 334001.0. ? YY = 48727.0

a)

120 100 80 60 40 20 400 500 200 300 100
c)

\\ \hat y = a + bx \\ a = \frac{\sum Y * \sum XX-\sum X*\sum XY}{n\sum X^2 -(\sum X)^2} \\ a = \frac{592.6*334001.0 - 1227.0*101296.1}{9*334001.0 - 1227.0*1227.0} \approx 49.0767 \\ b = \frac{n\sum XY -\sum X*\sum Y}{n\sum X^2 -(\sum X)^2} \\ b = \frac{9*101296.1 - 1227.0*592.6}{9*334001.0 - 1227.0*1227.0} \approx 0.123 \\ \hat y = 49.0767+0.123x

d)

\begin{tabular}{rrrrr} \hline X & Y & XY & XX & YY \\ \hline 63 & 17.7 & 1115.1 & 3969 & 313.29 \\ 146 & 81.7 & 11928.2 & 21316 & 6674.89 \\ 203 & 83.2 & 16889.6 & 41209 & 6922.24 \\ 33 & 69.2 & 2283.6 & 1089 & 4788.64 \\ 147 & 119 & 17493 & 21609 & 14161 \\ 36 & 63.7 & 2293.2 & 1296 & 4057.69 \\ 50 & 9.9 & 495 & 2500 & 98.01 \\ 62 & 55 & 3410 & 3844 & 3025 \\ \hline \end{tabular} \\ \sum X =740.0, \sum Y =499.4\\ \sum XY =55907.7, \sum XX =96832.0, \sum YY =40040.76
\\ \hat y = a + bx \\ a = \frac{\sum Y * \sum XX-\sum X*\sum XY}{n\sum X^2 -(\sum X)^2} \\ a = \frac{499.4*96832.0 - 740.0*55907.7}{8*96832.0 - 740.0*740.0} \approx 30.7686 \\ b = \frac{n\sum XY -\sum X*\sum Y}{n\sum X^2 -(\sum X)^2} \\ b = \frac{8*55907.7 - 740.0*499.4}{8*96832.0 - 740.0*740.0} \approx 0.3422 \\ \hat y = 30.7686+0.3422x